OpenStudy (anonymous):

Okay... I feel stupid for asking this.. but if I have 16 1/3 - 3 9/18ths... why can't I make them improper factions 48/3 - 54/18, then 288/18 - 54/18 to be 234/18 = 13... and that's wrong because the answer is supposed to be 12 and 5/6... aren't 288/18 and 16 1/3 the same number?

OpenStudy (asnaseer):

you made a mistake in turning them into improper fractions, e.g.:\[16\frac{1}{3}=16+\frac{1}{3}=\frac{16*3}{3}+\frac{1}{3}=\frac{48}{3}+\frac{1}{3}=\frac{49}{3}\]

OpenStudy (anonymous):

I understand that is why I got the answer wrong, but I don't understand WHY that makes the answer wrong. Shouldn't they be the same number if they're mixed numbers or improper fractions? So why should it matter what you do with them? They're the same number.

OpenStudy (asnaseer):

you should indeed get the same answer in either case, the mistake you made was in converting the mixed number to an improper fraction.

OpenStudy (anonymous):

So if it should be the same answer either way.. why isn't it? It comes very close to being the same, but it's slightly not the same.. so 16 1/3 and 288/18 are not actually the same number, just really similar?

OpenStudy (asnaseer):

you keep missing the point. your conversion from 16 1/3 to an improper fraction was wrong. it is not 48/3 - it is 49/3

OpenStudy (anonymous):

OHHH <3 Thank you!

OpenStudy (asnaseer):

yw - I'm glad I was (finally) able to help :)